OpenStudy (anonymous):

(3 pt) What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11?

OpenStudy (anonymous):

Any options?

OpenStudy (anonymous):

I go to K12 too btw xD

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

Im pretty sure its 11.

OpenStudy (anonymous):

No waitt

OpenStudy (anonymous):

|x| + |y + z| when x = –7, y = 8, and z = –11? |-7| + |8 + -11|

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

|-7| + |8 + -11| |8 + -11| = -3 -3 + -7 = -10

OpenStudy (anonymous):

those bars represent the absolute value, meaning the distance of the origin (regardless if negative or not). So, -7 would be changed to 7, and -3 (the sum of 8+-11) would be changed to 3. Adding 7 and 3 together, you get 10.

OpenStudy (anonymous):

Like i said but the total answer is |-10| because thats its absolute value.

OpenStudy (anonymous):

oh okay